Steve has 36 coins consisting of dimes and nickels. If the value of the coins is $3.10, then how many dimes does he have?
Thanks

Answer by stanbon(57203)

Steve has 36 coins consisting of dimes and nickels. If the value of the coins is $3.10, then how many dimes does he have?
---------------
Quantity Eq: d + n = 36
Value Eq: 10d +5n = 210 cents
------------------------------------
Multiply thru the Quantity equation by 5 to get:
5d + 5n = 5*36
---
Subtract that from the Value Eq. to get:
5d = 210-5*36
d = 42-36
d = 6 (# of dimes)
======================
Cheers,
Stan H.